an equilibrium mixture is analyzed and found to contain 0.94M N2 , 0.701M H2 and .35 M NH3 at room temp.

A. what is Kc for reaction N2(g) +3H2(g) -> 2NH3 (g) ?

B. write out the relationship between Kc and Kp, then calc Kp at 25C


please show all work
Answer #1

A)

Kc = [NH3]^2 / [N2][H2]^3

= (0.35)^2 / (0.94)*(0.701)^3

= 0.378

Answer: 0.378

B)

T= 25.0 oC

= (25.0+273) K

= 298 K

Δ n = number of gaseous molecule in product - number of gaseous molecule in reactant

Δ n = -2

use:

Kp= Kc (RT)^Δ n

Kp = 0.378*(0.08206*298.0)^(-2)

Kp = 6.321*10^-4

Answer: 6.32*10^-4